An arrow is shot straight up in the air at an initial speed of 15.0 m/s. After how much time is the arrow moving downward at a speed of 8.00 m/s?
1
Expert's answer
2020-10-08T13:27:49-0400

t time,

g the gravitational acceleration,

u the initial velocity

v the final velocity

"t=\frac{v-u}{a} = \frac{v-u}{g}=\frac{-8m/s - 15m/s}{-9.80665 m/s^2}=2.34535 seconds"

2.34535 sec